OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1/4x + 2 y = 4x - 1 y = 4x - 3 y = 1/4x - 3 y = 1/4x - 1

jimthompson5910 (jim_thompson5910):

the slope of y = 1/4x + 2 is 1/4

jimthompson5910 (jim_thompson5910):

any line parallel to y = 1/4x + 2 will also have a slope of 1/4 since parallel lines have equal slopes

jimthompson5910 (jim_thompson5910):

so we know that m = 1/4 the point (-4, -2) means that x = -4 and y = -2 plug this all into y = mx+b, then solve for b y = mx+b -2 = (1/4)(-4)+b ... ... ... b = ???

OpenStudy (anonymous):

b=-1 is that correct??

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

m = 1/4, b = -1 y = mx + b y = (1/4)x + (-1) y = 1/4x - 1

OpenStudy (anonymous):

THANK YOU!!!!!! You helped me a lot!!

jimthompson5910 (jim_thompson5910):

you're welcome, I'm glad I did